The sum of the squares of two consecutive odd integers is 34. Find the two integers.
Answer by checkley71(8403)

X^2+(X+2)^2=34
X^2+X^2+4X+4=34
2X^2+4X+4-34=0
2X^2+4X-30=0
2(X^2+2X-15)=0
2(X+5)(X-3)=0
X+5=0
X=-5 ANSWER.
X-3=0
X=3 ANSWER.
THEREFORE THE NUMBERS ARE (-3,3 & -5,5)